OpenStudy (firejay5):

What's the inverse function of f(x) = x - 5?

OpenStudy (anonymous):

i think y=3

OpenStudy (firejay5):

Show work on how you know it's the inverse.

OpenStudy (anonymous):

y= -3 replace y with x we get x = -3 this is the inverse

OpenStudy (firejay5):

that's what I put is there any work to show for it

OpenStudy (anonymous):

i think this work is enough which i did but i can do a graphical work for you

OpenStudy (firejay5):

like graphical meaning work it on a graph

OpenStudy (anonymous):

yes like if we gave a graph and its inverse they will always be symmetrical to the line y=x

OpenStudy (anonymous):

*have

OpenStudy (anonymous):

thus y = -3 and x = -3 are symmetrical to y = x

OpenStudy (firejay5):

I changed the question, so now it's f(x) = x - 5?

OpenStudy (firejay5):

@rizwan_uet think you help on this one

OpenStudy (anonymous):

f(x) = x - 5 y = x - 5 replace x with y and y with x x = y- 5 writting it in terms of y we get y = x+5 this is the required inverse

OpenStudy (firejay5):

can work be shown

OpenStudy (anonymous):

thats the work above

OpenStudy (firejay5):

so it's y = x - 5 & y = x + 5

OpenStudy (anonymous):

yes

OpenStudy (firejay5):

so there's really not much work you can show when working with inverse functions

OpenStudy (anonymous):

ofcourse its a few steps work

OpenStudy (firejay5):

can you show them to me

OpenStudy (anonymous):

com on please i showed the work to you. look it again f(x) = x - 5 y = x - 5 replace x with y and y with x x = y- 5 writting it in terms of x we get y = x+5 this is the required inverse

OpenStudy (firejay5):

I am sorry

OpenStudy (firejay5):

What would the inverse be of 2x + 3 all over 6

OpenStudy (anonymous):